write in polar form: 12+5i
Answer by fractalier(6550)

r = sqrt(x^2 + y^2)
theta = arctan(y/x) ...I use q for theta
here
r = sqrt(12^2 + 5^2) = sqrt(169) = 13
q = arctan(5/12)
so we have in polar form
(13, arctan(5/12))